### v3

# HDO ice detected toward an isolated low-mass protostar with JWST

Abstract
Water is detected in environments representing every stage of star and solar system formation, but its chemical evolution throughout these stages remains poorly constrained. Deuterium ratios offer a means of probing chemical links between water in different cosmic regions because of their sensitivity to physicochemical conditions. Here, we present the first detection of the 4.1 m HDO ice feature with JWST toward a low-mass protostar, L1527 IRS, which may eventually grow to a sun-like mass. We measure an ice HDO/H2O ratio of 4.410-3, where the reported error is dominated by uncertainties in continuum definition and ice band strengths. This fraction is similar to the gas HDO/H2O ratios measured in the warm (100 K) inner cores of other low-mass protostellar envelopes and protoplanetary disks found in comparably isolated star-forming regions. Such a similarity tentatively supports the assumption that water vapor detected in these regions is not significantly altered by gas-phase reactions following ice sublimation. It also supports the hypothesis that pre- and protostellar water ice is largely inherited in a chemically unaltered state by outer protoplanetary disks. However, the fraction is a factor of 4-10 times higher than the gas HDO/H2O ratios measured toward comets and low-mass protostars in clustered star-forming regions. This difference may be due to either gas-phase water reprocessing in protostellar envelopes and protoplanetary disks, or differences between prestellar conditions of isolated dense cores and the clustered star-forming regions that are more analogous to the environment in which our Sun formed.
1 Introduction
Water is a highly influential molecule that is essential to the formation and evolution of stars, planetary bodies, and life as we know it. It is also one of the most abundantly detected molecules in star-forming regions and has been observed at every evolutionary stage of star formation, from prestellar clouds to protostellar envelopes, protoplanetary disks, and solar system bodies such as our Earth, icy moons, and comets (Encrenaz, 2008; van Dishoeck et al., 2014, 2021).
In recent years, numerous works have investigated whether the water found in these various environments is chemically linked. Many of these studies examined water deuterium ratios because the sensitivity of deuterium isotopic exchange reactions to physicochemical conditions makes D/H ratios excellent tracers of the origins and evolutionary histories of molecules. In particular, much of the water detected in our solar system is D-enriched (e.g., Lécuyer et al. 1998; Waite et al. 2009; Alexander et al. 2012; Altwegg et al. 2015), in some cases by almost two orders of magnitude, with respect to the D/H observed in the diffuse interstellar medium (210-5, Prodanović et al. 2010). Because deuterium enrichment is efficient in low-temperature environments, these high water D/H ratios may be a remnant signature of water’s origins in the primitive ices of extremely cold and dense prestellar clouds from which stars and solar systems eventually form.
To confirm such a link between the water in our solar system and interstellar water ice, the water deuterium ratios (HDO/H2O) of ices in young star-forming regions must be measured. In theory, this value can be obtained by calculating HDO ice column densities via the 4.1 m O-D stretching mode of HDO. However, prior to the launch of the James Webb Space Telescope (JWST), reported detections of this feature (e.g., Teixeira et al. 1999) remained controversial due to the low sensitivity of previous IR telescopes (Dartois et al., 2003; Parise et al., 2003), hindering the direct quantification of reliable ice HDO/H2O ratios.
Interferometric (sub)millimeter gas-phase studies thus attempted to quantify interstellar ice HDO/H2O ratios indirectly instead by targeting the warm (100 K) inner regions of protostellar envelopes (“hot corinos”). Because it is often argued that the majority of the gas species detected in these environments originate from the thermal desorption of ices with little subsequent gas-phase chemical alteration (e.g., Requena-Torres et al. 2006; Coletta et al. 2020; Van Gelder et al. 2020; Nazari et al. 2021; Chen et al. 2023), their measured gas HDO/H2O ratios are assumed by some to reflect the ratios of their ice envelopes. Toward four protostars located in the clustered cloud complexes NGC 1333 and Ophiuchus, the measured gas HDO/H2O ratios range between 5-910-4 (Persson et al., 2014), very similar to the ratios measured in several comets in our solar system (e.g., Eberhardt et al. 1995; Villanueva et al. 2008; Altwegg et al. 2015). Because our solar system likely formed in a clustered star-forming region (Adams, 2010; Desch & Miret-Roig, 2024), this similarity has been used to assert the claim that cometary water ice was directly inherited from the prestellar cloud or protostellar envelope with minimal chemical processing in the protoplanetary disk in which comets formed (Cleeves et al., 2014). In contrast, the gas HDO/H2O ratios measured toward five distributed hot corinos found in isolated environments (i.e., they are not associated with star-forming cloud complexes) tend to be a factor of 2-10 higher (Jensen et al., 2019, 2021; Andreu et al., 2023; Tobin et al., 2023). Some of these studies suggested that such a difference in gas HDO/H2O ratios could be due to the lower radiation experienced by more isolated star-forming regions, which implies colder and/or longer prestellar stages than those in more clustered star-forming regions. It must be emphasized, however, that this reported dependence of the protostellar gas HDO/H2O on the local environment is based on a very limited sample size of sources, as discussed further in Section 3.3.
Nevertheless, it is possible that the gas-phase water detected toward hot corinos has experienced some degree of chemical processing in the gas phase (e.g., isotopic exchange with H2 gas), which would mean that the water accreted by bodies forming in protoplanetary disks could be substantially altered from its prestellar state. To ascertain that these gas HDO/H2O ratios are in fact not significantly altered by gas-phase processing and therefore are representative of ice HDO/H2O ratios, HDO ice must be quantified directly.
The advent of JWST, with its unprecedented sensitivity, offers an opportunity to revisit the search for HDO ice toward such star-forming regions. Recently, Slavicinska et al. (2024) demonstrated the capability of JWST to unambiguously detect HDO ice toward two massive protostars and subsequently calculated the ice HDO/H2O ratios of their envelopes. Here, we focus on the source L1527 IRS (also IRAS 04368+2557, hereafter L1527), an isolated low-mass class 0/I protostar (Tobin et al., 2013) found in the Taurus molecular cloud ( = 140 pc) that is deeply embedded within a dense 0.9 M⊙ envelope and has a 0.0075 M⊙, 75-125 AU-sized disk observed in the edge-on configuration (Aso et al., 2017; Kristensen et al., 2012; Tobin et al., 2013). The protostar currently has a mass around 0.3-0.5 M⊙ (Aso et al., 2017; van’t Hoff et al., 2023) but may eventually accrete enough material from its envelope to grow to a sun-like mass (Tobin et al., 2012).
A low-resolution (=/100) IR spectrum of the ices toward L1527 from 2.5-5.0 m was previously collected with the AKARI telescope. Aikawa et al. (2012) analyzed the ice features in this spectrum and tentatively assigned a weak feature observed close to the noise level at 4.13 m to HDO ice. While the 3 m H2O feature was saturated in this spectrum, an ice HDO/H2O ratio was calculated by estimating the H2O ice column density via fitting the wings of the 3 m feature. The resulting ratio, 2.010-2, is notably higher than the subsequently derived hot inner gas HDO/H2O ratios of other low-mass protostars (Persson et al., 2014; Jensen et al., 2019, 2021).
Here, we revisit the 4.1 m spectral region of L1527 with the JWST Near-InfraRed Spectrograph (NIRSpec), whose sensitivity, spatial resolution, and spectral resolution far exceed those of the AKARI facility. Furthermore, we use the JWST Mid-InfraRed Instrument (MIRI) to expand the observed spectrum of L1527 to 27.9 m, enabling the detection of the 12 m H2O ice libration feature. These new high-quality data allow us to determine a robust ice HDO/H2O ratio within an ice envelope of a low-mass protostar and compare this value with gas HDO/H2O ratios measured toward hot corinos. Such comparisons can serve to evaluate whether water in protostellar envelopes undergoes significant gas-phase reprocessing once it sublimates from ices, and whether such reprocessing affects the water that is eventually accreted by volatile-rich bodies formed in outer protoplanetary disks.
2 Methods
2.1 Observations
Medium-resolution (=/700-1300) IR spectra of L1527 covering the 4.1 m HDO ice band were collected with the NIRSpec Integral Field Unit (IFU) using the G395M grating as part of the cycle 1 GO program #1798 (PI John Tobin). NIRSpec IFU spectra collected with this grating provide continuous spectral coverage from 2.87-5.27 m, unlike spectra collected with the G395H grating, which have a detector gap that overlaps with the 4.1 m HDO ice feature. Additional low-resolution NIRSpec IFU PRISM (=/100) spectra from 0.6-5.3 m collected as part of the cycle 1 GO program #1960 (“JOYS+”, PI Ewine van Dishoeck), and MIRI Medium Resolution Spectrometer (MRS, =/1500-3500) spectra from 4.90-27.90 m collected as part of the cycle 1 GTO program #1290 (“JOYS”, PI Ewine van Dishoeck) gave continuous spectral coverage of L1527 from 0.6-27.9 m. These JWST data can be accessed in the Mikulski Archive for Space Telescopes (MAST) via https://doi.org/10.17909/jhkb-w045 (catalog doi: 10.17909/jhkb-w045).
The one-dimensional continuous 0.6-27.9 m spectrum analyzed in this work was extracted using 1” and 1.3” diameter apertures for the NIRSpec and MIRI datacubes, respectively (Figure 1). The aperture sizes differ to account for the PSF difference between NIRSpec and MIRI at the wavelength where their spectral coverages overlap (MIRI PSF at 5 m = 1.3 NIRSpec PSF at 5 m, Law et al. 2023; Federman et al. 2024). The spectra did not require scaling for the NIRSpec and MIRI fluxes to align in the overlapping 4.90-5.27 m wavelength range. The apertures were centered on positions that maximized the flux at 5 m (=4:39:53.929, =+26:03:09.617 in NIRSpec and =4:39:53.879, =+26:03:10.119 in MIRI) to correct for small (0.5”) differences in coordinate calibration between NIRSpec and MIRI. Further details regarding the extraction position and data reduction can be found in Appendix A.
2.2 Ice fitting
Although L1527 has a disk, its disk mass is 100 times less than the mass of its envelope – i.e., the source is still relatively young and very deeply embedded. Therefore, unlike in class II sources where the ice envelopes have largely dissipated and most of any observed ice absorptions are expected to originate from disks, we expect ices in the envelope to dominate the observed ice absorption toward L1527, and that the ice absorptions observed at short (NIRSpec) and long (MIRI) wavelengths probe similar spatial regions of the protostellar system. To confirm that this assumption is reasonable, we used the diagnostic technique of comparing ice bands of the same species at different wavelengths suggested by Pontoppidan et al. (2005). Specifically, we compared the profiles and optical depths of the 4.39 m 13CO2 and 3.53 m CH3OH bands in the NIRSpec data with those of the 15.2 m 12CO2 and 9.7 m CH3OH bands in the MIRI data and found that the degrees of thermal processing and column densities indicated by these bands were consistent with each other (see Appendix C). Thus, we opted to analyze the L1527 spectrum via classical techniques used to fit ice absorptions toward other class 0 and I sources (e.g., Gibb et al. 2004; Boogert et al. 2008) rather than the complex radiative transfer modeling employed to analyze ice absorptions toward class II sources (e.g, Sturm et al. 2023a; Bergner et al. 2024).
2.2.1 HDO and CH3OH ice fitting
The 4 m spectral region where the strongest HDO ice feature is found can also include contributions from CH3OH and S-bearing ice species. This work used the procedures and laboratory data described in Slavicinska et al. (2024) to deconvolve these blended features and quantify the HDO and CH3OH ice observed toward L1527 via the NIRSpec G395M data.
To briefly summarize this procedure, the CH3OH ice column density was first obtained by fitting the 3.53 m feature (isolated via a third-order polynomial local continuum subtraction, see Figure B.4) with CH3OH ice laboratory spectra and a Gaussian representing NH3 hydrates, although NH4+ salts can also have similar broad absorption feature around the same wavelength (Schutte & Khanna, 2003). More details on the CH3OH ice fitting at 3.53 m can be found in Appendix B.3. The resulting CH3OH ice column density, calculated using a band strength of 4.8610-18 cm molec-1 for the 3.53 m band from Luna et al. (2018), was then used to constrain the expected contribution of the CH3OH combination modes to the 3.7-4.2 m spectral region (isolated via a fourth-order polynomial local continuum subtraction, see Figure B.1). The remaining spectral excess between 3.7-4.2 m was subsequently fit with a combination of a Gaussian representing SH-bearing species and smoothed laboratory spectra of both amorphous and crystalline HDO ice. Band strengths of 6.910-17 and 6.810-17 cm molec-1 from Table A2 in Slavicinska et al. (2024) corresponding to the temperatures of the best-fitting amorphous and crystalline HDO profiles, respectively, were used to calculate the total HDO ice column density. These corresponding temperatures should not be interpreted as directly representative of the ice temperatures, as this two-component fit is only a first-order approximation of the continuous thermal gradient that is expected along the line of sight. Details on the uncertainty calculations for the HDO and CH3OH ice column density can be found in Appendix D.
2.2.2 H2O ice fitting
H2O ice, the most abundant observable ice toward most protostars, has mid-IR absorption features at approximately 3, 4.5, 6, and 12 m that vary significantly in strength (Mastrapa et al., 2009). The 3 m stretching feature is preferentially used to quantify H2O ice whenever possible because out of the four main H2O mid-IR features, it is the strongest and the least blended with other ice or dust bands. However, as a result of its strength and the relatively high abundance of water ice, it is frequently saturated toward the most embedded protostellar sight-lines, as is the case for L1527. Two alternative methods can be used to derive H2O ice column densities when this is the case: 1) fitting one or both of the wings of the 3 m feature (e.g., Aikawa et al. 2012), or 2) fitting the 12 m libration feature (e.g., Boogert et al. 2008; Chen et al. 2024). The first method is complicated by the fact that the wings of the 3 m band shift with temperature, grain size, and chemical environment (Smith et al., 1989; Dartois & d’Hendecourt, 2001), so H2O ice column densities derived from this method are highly dependent on the physical and chemical parameters of the template H2O ice spectra used in the fit, as discussed further in Appendix E. The second method requires full mid-IR spectral coverage due to the broad profile of the 12 m feature as well as its blending with similarly broad neighboring silicate features at 10 and 18 m.
In this work, we derived the H2O ice column density by fitting the 12 m band. We used the observed optical depths at 2.8, 4.5, and 6 m as additional constraints (i.e., the blue wing of the 3 m band and the 4.5 and 6 m features should not be vastly overfit with the model). Details on the continuum fitting used to extract the spectrum into the optical depth scale and the silicate subtraction used to isolate the 12 m band can be found in Appendix B. A set of H2O ice spectral models with a range of temperatures and grain sizes were generated via the DHS model for irregular grains (Min et al., 2005) with the OPTOOL code (Dominik et al., 2021), using water ice optical constants from (Mastrapa et al., 2009). The full continuum- and silicate-subtracted spectrum was then fit with various combinations of these models until a combination that met all of the fitting criteria (sufficiently fits the 12 m band without vastly overfitting the blue wing of the 3 m band and the 4.5 and 6 m bands) was found. The final model used a power-law size distribution (a) a-3.5 with grain sizes ranging from 0.1-1.0 m and included amorphous H2O ice at 40 K, amorphous H2O ice at 80 K, and crystalline H2O ice at 80 K. Band strengths from Mastrapa et al. (2009) corresponding to these respective ice temperatures and phases (2.310-17, 3.010-17, and 3.310-17 cm molec-1) were then used to calculate the total H2O ice column density by summing the contribution to the full 12 m band from each of these fit components. As with the HDO, the temperatures used in this model should not be interpreted as extracted temperatures of the H2O ice, as they are only a rough approximation of the expected continuous thermal gradient along the line of sight. Additionally, these fits use only pure H2O ice laboratory data and thus do not account for possible changes in the peak profiles due to chemical effects (e.g., mixing with other ice species like CO2, NH3/NH4+, CH3OH, etc.), which can be degenerate with temperature effects. Details on the uncertainty calculations for the H2O ice column density can be found in Appendix D.
3 Results and discussion
3.1 HDO ice
Figure 2 presents the 3.7-4.2 m region of the L1527 spectrum in optical depth units after the local continuum subtraction (see Appendix B). The CH3OH column density from the 3.53 m fit (see Appendix B.3) applied to fit this spectral region leaves small (0.01-0.02 ) and large (0.14 ) excess absorptions between 3.9-4.0 and 4.0-4.2 m, respectively. The large excess between 4.0-4.2 m is observed throughout the entire datacube, including areas both with and without scattering wings around the 4.27 m CO2 band (see Figure B.1).
The absorption excess between 4.0-4.2 m presents an asymmetric feature with a profile very similar to that observed by Slavicinska et al. (2024) toward higher mass sources, although with much higher optical depth and improved S/N. Given an RMS value of 0.004 determined from the relatively featureless 3.71-3.77 m range, the feature is detected at almost 40, far exceeding the S/N of the L1527 HDO ice detection reported by Aikawa et al. (2012). As with the massive protostellar spectra presented in Slavicinska et al. (2024), the feature is fit well using a combination of amorphous and crystalline HDO ice profiles that correspond to temperatures of 100 and 70 K, respectively, in laboratory 0.4% HDO:H2O ice mixtures. Such a combination of amorphous and crystalline components approximates to a low degree a thermal gradient within the ice envelope, which is expected along the line of sight, and thus should not be interpreted as the extracted physical ice temperatures. The presence of both amorphous and crystalline HDO is consistent with additional spectral evidence for thermal processing toward this source, such as the double-peaked 4.39 m 13CO2 stretching and 15.2 m 12CO2 bending features (Pontoppidan et al., 2008; Brunken et al., 2024b), and indicates that the most heated regions have experienced high enough temperatures to crystallize water ice. The HDO ice column density calculated from these fits, 9.21016 cm-2 (see Table 1), is consistent with the L1527 HDO ice column density of 9.61016 cm-2 reported from AKARI data by Aikawa et al. (2012) despite the differences in the data quality and fitting methods (further discussion in Appendix E.1).
The excess absorption between 3.9-4.0 m is fit well with a Gaussian with a peak position of 3.9 m and FWHM of 0.1 m, which is similar to the profile of the S-H stretching mode of SH-bearing species like H2S and NH4SH in water-rich ices (Yarnall & Hudson, 2022; Slavicinska et al., 2025). The fit presented in Figure 2 results in column density upper limits of 0.3% with respect to H2O if assuming the only species contributing to the excess is H2S and 13% with respect to H2O if assuming the only species contributing to the excess is NH4SH, but these upper limits are extremely dependent on the continuum choice, resulting in estimated uncertainties on the order of 300% and 70%, respectively (see Appendix D).
| Ice | Feature | Optical depth | Column density | Abundance |
| m | 1017 cm-2 | w.r.t. H2O | ||
| H2O | 12 | 2.0 | 209 | 1 |
| HDO | 4.1 | 0.14aaOptical depth of the entire amorphous and crystalline HDO absorption complex at 4.14 m. The optical depths of the individual fitted amorphous and crystalline components are 0.05 and 0.101, respectively, with the asymmetric errors reflecting the differences in the optical depth of the fit components when using the “upper and lower limit” local continua shown in Appendix D. | 0.92 | 4.410-3 |
| CH3OH | 3.53 | 0.58bbOptical depth of the entire CH3OH+NH3 absorption complex at 3.53 m. | 22 | 0.11 |
3.2 H2O ice
The profile of the isolated 12 m H2O libration band is fit with a combination of grain-shape corrected models of amorphous H2O ice at 40 K, amorphous H2O ice at 80 K, and crystalline H2O ice at 80 K (Figure 3). Such a combination of amorphous and crystalline components again approximates a thermal gradient within the ice envelope. This fit at 12 m is consistent with the optical depth of the 4.5 m combination mode and does not result in any overfitting of the 6 m feature (which consists of the H2O ice bending mode blended with other ice features, Boogert et al. 2008). However, the blue wing of the 3 m band is slightly overfit, while the red wings of the 3 and 12 m bands are underfit. It was not possible to obtain a satisfactory fit to these wings while still sufficiently fitting the 12 m feature using only grain shape-corrected pure H2O models. Including laboratory spectra of H2O mixed with other major ice components (e.g., CO2, NH3, ammonium salts) may result in improvements to the fits in these wings, but we did not include such components in our fit for simplicity’s sake, as deducing the chemical environment of H2O ice is outside the scope of this work, and the inclusion of such components likely would not affect the final calculated H2O ice column density outside of the reported uncertainties (see Appendix D) given that the band strength of the H2O libration mode is not strongly affected by mixing with other common interstellar ice constituents (Kitta & Krätschmer, 1983).
The H2O ice column density obtained from the fit, 2.091019 cm-2 (see Table 1), is a factor of 4 greater than the H2O ice column density measured from the AKARI spectra (Aikawa et al., 2012), which was likely underestimated due to the availability of only the wings of the 3 m band for fitting and a single template H2O ice spectrum at a very cold temperature (10 K) being used in the fit (see Appendix E.2). A higher H2O ice column density, on the order of 1019 molec cm-2, is expected toward this source given its massive envelope and very deep silicate feature (Figure E.1). In fact, the high water column density and optical depth of the 9.7 m silicate feature (4.2) of L1527 are typical of other deeply embedded low-mass Class 0 protostars now being probed by JWST (e.g., Chen et al. 2024).
While the 12 m H2O feature serves as a satisfactory means of measuring the total H2O ice column density, it cannot reliably indicate ice physical properties in the way that the sharp, asymmetric profile of the 4.1 m HDO band indicates crystallinity (Section 3.1) due to the degenerate effects of temperature, grain shape and size, chemical environment, and the choices in continuum and subtracted silicate template on its broad profile (e.g., Boogert et al. 2008; Mastrapa et al. 2009). This means that the temperatures of the H2O ice data used in the H2O fit should not be interpreted as representative of the actual temperature of the observed H2O ice, as possible effects of different chemical environments on the profile of the 12 m band were not considered here (see Section 2.2.2). The profile change of the 12 m feature upon crystallization is also not as dramatic compared to the very evident sharpening of the 3 m H2O stretching mode, which is saturated toward this source. As mentioned in Section 2.2.2, physicochemical effects on the wings of the 3 m feature can be degenerate as well, so these wings are also unreliable tracers of physical properties when the 3 m feature is saturated. For this reason, no secure conclusions can be drawn about the temperature or morphology of H2O ice toward this source from this fit.
However, the total L1527 water ice abundance with respect to the hydrogen column density can provide other valuable information regarding the thermal processing history of the ice envelope. This value can be approximated using the relation between silicate optical depth and in the dense cores of Lupus (Boogert et al., 2013). The resulting inferred H2O ice abundance of L1527 is 410-5, on par with the higher end of H2O ice abundance values measured toward other low-mass protostars (Figure E.1). Such an abundance demonstrates that, although temperatures high enough to crystallize water ice must have been reached in at least some parts of the envelope of L1527 based on the 4.1 m HDO ice band profile, a large portion of the envelope was likely not heated to high enough temperatures to desorb a significant portion of the H2O ice (100 K). The relatively blue peak position of the L1527 6.85 m NH4+ feature, which is more redshifted toward sources that have lost most of their H2O ice, is also consistent with this moderate processing scenario (Slavicinska et al., 2025).
3.3 HDO/H2O ratio
The column densities of HDO and H2O ice found here result in an ice HDO/H2O ratio of 4.410-3. A comparison of this ratio with HDO/H2O values and upper limits previously measured in protostellar envelopes as well as volatile-rich solar system objects is presented in Figure 4.
Within the reported uncertainties, our value is consistent with the range of HDO/H2O gas values that have been observed toward isolated low-mass protostars (1.80.4-6.31.510-3) in the hot inner regions of Class 0 and I protostellar envelopes (Jensen et al., 2019, 2021; Andreu et al., 2023). This similarity supports the assumption that the gas HDO/H2O ratios measured in the hot regions of protostars with sun-like masses are indeed representative of their ice HDO/H2O ratios. Nevertheless, given the nearly half order of magnitude scatter observed in the gas HDO/H2O ratios of isolated low-mass protostars, it is clear that parameters such as specific local conditions and evolutionary timescales may affect these ratios as well, so future work comparing gas and ice HDO/H2O ratios toward the same protostar will be crucial for constraining the exact extent of gas-phase water reprocessing upon the sublimation of water ice in the envelopes of these objects.
Tobin et al. (2023) recently showed that the gas HDO/H2O ratio of 2.260.6310-3 measured within the extended snowline of a disk of an isolated Class I protostar that recently experienced an accretion burst, V883 Ori, is generally consistent with the gas ratios measured in the hot inner regions of isolated Class 0 low-mass protostars. Such a consistency serves as evidence that the ices formed early in the star formation process, in prestellar clouds and protostellar envelopes, may be largely inherited by the outer regions of protoplanetary disks (and, by extension, the planetary bodies that form within them) of sun-like protostars without substantial chemical alteration via gas-phase reactions. Given that L1527 is an isolated low-mass protostar classified somewhere between the Class 0 and I stages (Tobin et al., 2013), the similarity of its ice HDO/H2O ratio to gas HDO/H2O ratios measured toward both Class 0 and Class I isolated sources, including V883 Ori, further corroborates this scenario.
The ice HDO/H2O ratio of L1527 being consistent with the gas HDO/H2O ratios of other isolated low-mass protostars means that it is also higher by a factor of 4-7 than the gas HDO/H2O ratios measured in the hot inner regions of clustered Class 0 protostars and up to an order of magnitude higher than the HDO/H2O ratios measured in comets (see Tables F.1 and F.2). As mentioned previously, it has been suggested that water formed in isolated star-forming regions has higher D/H ratios than that formed in clustered regions because of either longer or colder prestellar stages, during which the majority of the ices in protostellar envelopes form (Jensen et al., 2019, 2021). If this hypothesis is correct and local conditions greatly influence deuterium ratios, then our L1527 ice HDO/H2O ratio is not fully comparable to cometary HDO/H2O ratios because our Sun likely formed in a highly clustered and irradiated (and thus warmer) star-forming region (Adams, 2010; Desch & Miret-Roig, 2024). However, more samples are needed to ensure that this trend of higher HDO/H2O ratios in more isolated protostars is statistically robust, as the current sample of protostellar gas HDO/H2O values is still relatively small and consists of sources scattered throughout multiple molecular clouds in different galactic locations. The four currently published clustered protostellar gas HDO/H2O values were measured toward the Ophiuchus and Perseus NGC 1333 molecular clouds, while the five isolated gas values were measured toward less populated regions in Orion and Taurus and three different isolated Bok globules. This means that any comparison of their values does not control for differences that could arise due to sampling from different star-forming regions. For this reason, larger sample sizes of both gas and ice HDO/H2O measurements, particularly of protostars in the same molecular cloud but with different proximities to other stars, are crucial.
Furthermore, the ice HDO/H2O ratio of L1527 is very similar to the ice HDO/H2O ratios measured recently with NIRSpec IFU G395M toward more massive protostars (2.61.4 and 4.62.2 10-3, Slavicinska et al. 2024). HOPS 370 in particular is located north of the extended Orion Nebula Cluster, a highly active star-forming region, so its similar ice HDO/H2O ratio to L1527 is inconsistent with the above hypothesis. It is also inconsistent with the trend observed in the deuterium ratio of CH3OH gas in protostars, which correlates negatively with stellar mass (i.e., more massive protostars tend to have lower CH2DOH/CH3OH ratios, van Gelder et al. 2022). If isotopic ratios of species that formed primarily in the solid state do in fact reflect prestellar cloud conditions as has been previously suggested (Jensen et al., 2019; van Gelder et al., 2022), the similarity between HDO/H2O ratios in low- and high-mass protostars, both in the ice as noted here and in some gas measurements as discussed previously in Slavicinska et al. (2024), would imply similar prestellar conditions between isolated low-mass protostars and some massive protostars. However, the large scatter in the massive protostellar gas HDO/H2O ratios and the small ice HDO/H2O sample size in both isolated low-mass and massive protostars precludes drawing more general statements regarding the prestellar conditions of low-mass and massive protostars as a whole.
4 Conclusions
We report the first JWST detection and quantification of HDO ice toward a low-mass protostar, L1527 IRS. The detection and quantification were performed using the 4.1 m O-D stretching mode observed with the NIRSpec IFU instrument. Due to JWST’s unprecedented sensitivity and spectral resolution, the detection exceeds the noise level by almost 40, and the feature is spectrally resolved enough to reveal an asymmetric profile characteristic of both amorphous and crystalline HDO ice in the envelope of L1527. We subsequently quantified the ice HDO/H2O ratio toward this source using the 12 m H2O ice libration mode in the MIRI range. Our main conclusions can be summarized as follows:
1.
We measured an ice HDO/H2O ratio toward L1527 of 4.410-3, where the reported uncertainties are dominated by uncertainties in continuum definition and laboratory ice band strengths rather than observational uncertainties.
2.
The L1527 ice HDO/H2O ratio is similar to the gas HDO/H2O ratios measured interferometrically toward the hot inner regions of other isolated low-mass protostars and within the snowline of the young protoplanetary disk V883 Ori. This similarity is in general agreement with the assumption that gas-phase molecular abundances in hot corinos and young disks are representative of those in protostellar ices, with minimal alteration caused by chemical processing by gas-phase reactions following ice sublimation.
3.
The ice HDO/H2O ratio of L1527 is a factor of 4-7 times greater than the gas HDO/H2O ratios measured toward Class 0 clustered low-mass protostars and up to an order of magnitude greater than those measured in comets. This difference may be due to the water in these objects having undergone gas-phase reprocessing at some stage, or because of differences in the local star-forming environments being reflected in the HDO/H2O ratio of ices formed during the prestellar stages. If the latter scenario is the case, then the ice HDO/H2O ratio measured toward an isolated object like L1527 would be expected to exceed the ratios measured in comets in our solar system because our Sun likely formed in a more clustered environment.
Moving forward, a comparison of the gas and ice HDO/H2O ratios measured toward the same source will be essential to conclusively determine the extent of gas-phase chemical alteration of water in protostellar envelopes and protoplanetary disks given the large scatter in gas HDO/H2O values measured in low-mass protostars. Additionally, measurements of HDO ice toward a larger sample of low-mass protostars in both clustered and isolated star-forming regions are needed to fully characterize the effects of local environmental conditions on ice HDO/H2O ratios in protostellar envelopes and to conclusively evaluate the plausible link between protostellar and cometary water ice. The unprecedented sensitivity and spectral resolution of JWST opens the door for such measurements in the future.
5 Acknowedgments
K.S. thanks Brian Ferrari and Gary Huss for helpful scientific discussions. The authors thank an anonymous reviewer for their insightful comments that helped to improve this paper. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1798, #1290, and #1960. All the data used in this paper can be found in MAST: http://dx.doi.org/10.17909/jhkb-w045 (catalog 10.17909/jhkb-w045). The following National and International Funding Agencies funded and supported the MIRI development: NASA; ESA; Belgian Science Policy Office (BELSPO); Centre Nationale d’Etudes Spatiales (CNES); Danish National Space Centre; Deutsches Zentrum fur Luft- und Raumfahrt (DLR); Enterprise Ireland; Ministerio De Economiá y Competividad; Netherlands Research School for Astronomy (NOVA); Netherlands Organisation for Scientific Research (NWO); Science and Technology Facilities Council; Swiss Space Office; Swedish National Space Agency; and UK Space Agency. Astrochemistry at Leiden is supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101019751 MOLDISK), the Netherlands Research School for Astronomy (NOVA), and the Danish National Research Foundation through the Center of Excellence “InterCat” (Grant agreement no.: DNRF150). A.C.G. acknowledges support from PRIN-MUR 2022 20228JPA3A “The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 “NIR-dark Accretion Outbursts in Massive Young stellar objects (NAOMY)” and Large Grant INAF 2022 “YSOs Outflows, Disks and Accretion: towards a global framework for the evolution of planet forming systems (YODA)”.
Appendix A Observations
The extraction coordinates are centered on the upper layer of the L1527 disk seen in scattered light, slightly off the source (which lies in the dark area between the two scattered light lobes visible in Figure 1) to maximize S/N of the ice features. However, besides S/N differences, there is little variation in the profiles and relative optical depths of the ice features in the spectrum extracted from these coordinates and the spectrum extracted from the source center, and in fact, there are no noticeable spatial variations in the ice band profiles throughout most of the datacube.
The NIRSpec data were reduced via the JWST pipeline v1.16.0 using the Calibration Reference Data System 154 (CRDS) context file jwst_1298.pmap. A wavelength offset of -0.004 m was applied to the PRISM data to correct for what appears to be a discrepancy in wavelength calibration between PRISM and the G395M/H gratings observed in peak positions of both gas and ice lines. The MIRI data reduction is described in Rangaswamy et al. (in prep.).
Appendix B Continuum, silicate, and CH3OH ice fitting
B.1 Local continuum at 3.7-4.2 m
The top panel of Figure B.1 shows the continuum applied to extract the optical depth spectrum of the 3.7-4.2 m region that was used to calculate the HDO ice column density (Figure 2). The local continuum fitting of this region is not trivial in this spectrum because the continuum is distorted by light scattering at the wings of the 4.27 m CO2 ice band, resulting in a break between the respective continua on the blue and red sides of the feature (e.g., Dartois et al. 2022).
However, in other spatial regions within the datacube, data can be extracted where the CO2 scattering wings are minimal, and such a break in the continuum around the feature does not occur. The middle panel of Figure B.1 shows a spectrum extracted from such a region located east of the source. For this spectrum, the local continuum fitting at 3.7-4.2 m is more straightforward as it can be defined so that it follows the flux on both sides of the 4.27 m CO2 band. This off-source spectrum can therefore be used to guide the local continuum fit to the higher S/N spectrum extracted from the region closer to the center of the source, where the continuum around the 4.27 m band is disrupted by scattering.
The extracted spectra in optical depth resulting from these two local continuum fittings are shown in the bottom plot in Figure B.1. Although the local continua of the two spectra differ in shape, the profiles and relative optical depths of the extracted ice features are consistent with each other. The profiles and relative optical depths of ice features in spectral regions where local continuum fitting is more trivial because there are no spectrally adjacent strong features distorting the continuum (e.g., CH3OH at 3.5 m, 13CO2 at 4.4 m, and OCN- and CO at 4.6-4.7 m) are also consistent, which shows that the ices observed along these two lines of sight have similar chemical compositions and have experienced similar degrees of thermal processing. However, the absolute optical depths of the ice features observed near the source center are on average a factor of 1.26 higher than those of the features in the offset position, which may be due to the ice envelope toward the offset position being thinner. For this reason, the offset spectrum in the bottom panel of Figure B.1 is scaled by a factor of 1.26. The consistency between the extracted HDO absorption profiles provides confidence that the local continuum fit to the flux extracted from the central position sufficiently models the continuum distortion caused by the scattering wings of the 4.27 m CO2 ice band.
B.2 Global continuum and silicate subtraction
To extract the full combined NIRSpec+MIRI spectrum into the optical depth scale, we defined two separate but connecting global continua in the near-IR (2.5-5 m) and mid-IR (5-28 m) ranges by fitting low-order polynomials to the data on a log(Fλ) versus log() scale, similar to the procedure described in Boogert et al. (2008). The resulting global continuum fit to the 2.5-27.9 m spectral range is shown in Figure B.2. Similarly curved SEDs have been used to subtract the global continuum from other protostellar spectra (e.g., see the CrA IRAS 32, IRAS 03235+3004, and CRBR 2422.8-3423 global continua in Figure 1 of Boogert et al. 2008) and can be reproduced by radiative transfer models of protostellar systems (Crapsi et al., 2008). In the case of our global continuum for L1527, the polynomial continuum fit can also be reproduced well with two blackbody curves, one at 450 K and another at 100 K, roughly simulating scattered light and thermal emission from different regions of the L1527 system.
We then isolated the 12 m feature by subtracting a synthetic silicate spectrum generated for small spherical grains via the optical constants of amorphous olivine (MgFeSiO4) and pyroxene (Mg0.5Fe0.5SiO3) from Dorschner et al. (1995), as done previously for other dense and icy sightlines (Boogert et al., 2011). The fit used for the silicate subtraction is shown in Figure B.3.
B.3 Local continuum at 3.3-3.7 m and CH3OH ice fit
The fit to the 3.3-3.7 m region used to calculate the CH3OH ice column density is shown in Figure B.5. Despite the obvious presence of thermally processed ices along the line of sight, the CH3OH feature at 3.53 m is best fit only with laboratory spectra of cold (40 K) CH3OH. Additionally, a combination of pure CH3OH and CH3OH in a water-dominated ice (in this case CH3OH:H2O 1:5) is needed to fit the feature. These fitting parameters are very similar to those used in Slavicinska et al. (2024) toward two massive protostars with thermally processed lines of sight. The resulting CH3OH ice column density is 2.21018 cm-2. The abundance of CH3OH ice with respect to H2O ice, 11%, is close to the upper quartile value of CH3OH ice abundances toward the sample of mostly Class I low-mass protostars observed before JWST (Boogert et al., 2015).
Appendix C Comparing NIRSpec and MIRI bands of the same ice species
Comparing the profiles and optical depths of bands found at different wavelengths but belonging to the same ice species can serve as a tool to evaluate whether disk radiative transfer effects need to be considered toward a source, as the ice column densities and chemical environments extracted from such bands of common species can differ in the same line of sight due to differences in photon paths through the disk at different wavelengths (Pontoppidan et al., 2005; Sturm et al., 2023b). Such differences can indicate that the observed ice absorptions originate from ices located in different regions of the protostellar system (i.e., the short and long wavelengths do not probe the same spatial regions). Here we compare the 13CO2 stretching mode at 4.39 m and the 12CO2 bending mode at 15.2 m as well as the CH3OH C-H stretching mode at 3.53 m and the CH3OH C-O stretching mode at 9.7 m to investigate if the ice column densities and physicochemical environments extracted from these profiles at different wavelengths are consistent with each other, as would be expected for ice absorptions that originate from an isotropic envelope made up of small (1 m) grains, where the same complex radiative transfer effects are not at play.
Figure C.1 shows the comparison of these bands. First, both the 4.39 m 13CO2 feature and the 15.2 m 12CO2 feature have double-peaked profiles indicative of ice thermal processing (see Boogert et al. 2000; Pontoppidan et al. 2008; Brunken et al. 2024b). This consistency demonstrates that the degrees of thermal processing experienced by the ices probed at short and long wavelengths are similar. Furthermore, the 12C/13C ratio derived from these two bands by Brunken et al. (2024a) for L1527, 105, is consistent with the range of 12C/13C ratios derived for less evolved class 0 sources with split CO2 features (100-120). Although the 4.27 m 12CO2 stretching feature is saturated toward L1527 and the narrow 2.70 m 12CO2 combination mode is not resolved well enough in the low-resolution PRISM spectrum to derive a 12CO2 column density from the shorter wavelengths, Brunken et al. (2024a) found that for similar sources with similar 12C/13C ratios and coherent 2.70, 4.27, and 15.2 m bands, the 12CO2 column densities derived from NIRSpec and MIRI data are remarkably consistent (typically within 5-20% error).
The CH3OH ice column density derived from fitting the 9.7 m band using the same combination of pure CH3OH and CH3OH:H2O spectra used to fit the 3.5 m absorption complex results in a CH3OH column density of 3.10.61018 molec cm-2 using a band strength of 1.5610-17 cm molec-1 from Luna et al. (2018). This CH3OH column density is consistent with the CH3OH column density derived from the 3.53 m band within the reported uncertainties that take into account both uncertainties in the continuum subtraction and band strengths. The difference between the two values, 33%, is similar to the difference between CH3OH column densities from the 3.53 and 9.7 m bands reported for observations of dense clouds (McClure et al., 2023).
These consistencies in the ice morphologies and column densities between ice bands of the same species probed at different wavelengths support the notion that the short and long wavelengths probe similar spatial regions in the L1527 protostellar system and provide confidence that the ice absorptions observed toward L1527 can be analyzed using classical ice fitting techniques.
Appendix D Ice fitting uncertainties
The reported uncertainties on the HDO, H2O, and CH3OH ice column densities were determined here using methods similar to those described in Slavicinska et al. (2024). Two main sources of uncertainty were considered: uncertainty of the local/global continuum choice, and uncertainty of the laboratory band strengths used to calculate the ice column densities. Compared to these two sources of uncertainty, other sources of uncertainty like observational noise and fitting covariances are small and are therefore treated here as negligible.
We quantified the uncertainties on the column densities from continuum uncertainty by defining the lower and upper limits of what we consider to be a “reasonable” continuum (top panels of Figures D.1, D.2, and D.3) and then using the same methods described in Section 2 to derive lower and upper limits of HDO, H2O, and CH3OH column densities by fitting the optical depth spectra extracted via these continua (bottom panels of Figures D.1, D.2, and D.3). In the case of HDO, we also ensured that a similar extracted spectral profile could be achieved with a “reasonable” local continuum fit to the spectrum extracted from the off-center position discussed in Section B.1 (second and third rows in Figure D.1). These column density lower and upper limits serve as estimates of the column density uncertainties from uncertainties in continuum fitting. They are asymmetric and, to an extent, subjective depending on what one considers to be a “reasonable” continuum.
The reported laboratory band strengths of the HDO and H2O bands used here to calculate the ice column densities can vary significantly with temperature, particularly for the amorphous ices. Because we do not attempt to model the temperature gradient with our fits, we treat this variation of band strength with temperature as an additional source of uncertainty in the ice column densities. We again calculate lower and upper limits of HDO and H2O ice column densities, this time by using the respective highest and lowest band strength measured for a given phase of the ice. The resulting uncertainties are again asymmetric. For CH3OH, temperature and chemical environment changes can change the band strength of the 3.53 m feature by a maximum of 20% (Kerkhof et al., 1999; Luna et al., 2018), and Luna et al. (2018) report experimental uncertainties of 20%. For all three ice species, the band strength and continuum uncertainties are both considered systematic and are thus added linearly to provide the total ice column density uncertainties.
Although SH-bearing ice species are not the focus of this work, we report in Section 3.1 H2S and NH4SH upper limits (0.3 and 13% w.r.t. H2O, respectively) calculated from the Gaussian fit to the excess at 3.9 m, in the case that these upper limits may be useful for other studies of sulfur in star-forming regions or interstellar ices. However, it is evident from Figure D.1 that the optical depth of this Gaussian can vary dramatically with the continuum choice, resulting in high uncertainties in these reported upper limits. Depending on the continuum, the H2S upper limits can range from 0.07-2%. For NH4SH, the upper limit is 4% in the case of the lower limit continuum choice shown in the bottom left panel of Figure D.1. In the case of the upper limit continuum choice in the bottom right panel, the 3.9 m Gaussian profile is too broad to be attributed exclusively with NH4SH. Furthermore, the L1527 NH4+ column density derived from the 6.85 m feature (Slavicinska et al., 2025) provides an additional constraint that the NH4SH upper limit cannot exceed that derived from the fit presented in Figure 2 (13%).
Appendix E Comparison with AKARI results
E.1 HDO ice
While Aikawa et al. (2012) cautioned that their HDO column densities obtained from AKARI data toward L1527 “should be interpreted with caution” given the low signal-to-noise (“small bumps and hollows that deviate from the HDO feature”), the HDO ice column density measured here is similar to their value. Although the AKARI data were not spectrally resolved enough to reveal the asymmetric profile of the 4.1 m band and were thus only fit with a single Gaussian, Aikawa et al. (2012) noted that the profile of their best-fitting Gaussian was narrower and peaked at longer wavelengths than expected for cold and amorphous HDO ice and thus stated that the feature resembles HDO ice that has been annealed. While this is in contradiction with their fitting of 10 K amorphous H2O ice to the 3 m band wings, it is in agreement with our analysis that includes a prominent crystalline component of the HDO band.
E.2 H2O ice
Our H2O ice column density, 2.091019 cm-2, is over a factor of 4 greater than the H2O ice column density measured toward L1527 using AKARI spectra by Aikawa et al. (2012), 4.71.31018 cm-2. This is because Aikawa et al. (2012) derived their H2O ice column densities by fitting the wings of the 3 m feature with pure, 10 K H2O ice spectra. Because the wings of this feature can redshift significantly with grain effects and higher temperatures, it is easy to over- or underfit this feature when fitting only its wings. Aikawa et al. (2012) state that “the H2O ice column density cannot be significantly larger than we estimated” because their fit to the 3 m band wings is also consistent with their extracted optical depth of the 4.5 m combination mode, but the optical depth of this extremely broad feature can change significantly with even small changes in the fitted global continuum. In past literature, the continuum from 2-5 m has frequently been defined using datapoints around 4 m (e.g., Gibb et al. 2004) despite the fact that numerous weak ice features, including the broad blue wing of the H2O combination mode, have absorptions in this spectral region. Such fitting can result in an underestimation of the H2O ice combination mode optical depth.
Further supporting our significantly larger L1527 H2O ice column density is a comparison of the 9.7 m silicate feature optical depth and the H2O ice column density observed toward L1527 versus those of other low-mass protostars (Figure E.1). In general, the optical depth of the 9.7 m feature strongly correlates with the observed H2O column density toward these objects, and based on the high 9.7 m feature optical depth observed toward L1527 ( 4.2), its H2O ice column density is indeed expected to be 21019 cm-2.
Appendix F Literature HDO/H2O values
| Object | Phase | HDO/H2O | Ref. |
| 10-3 | |||
| Carbonaceous Chondrites | |||
| CI chondrites | - | 0.129-0.195 | 1 |
| CM chondrites | - | 0.1730.0072 | 1 |
| CR chondrites | - | 0.342 | 1 |
| CO chondrites | - | 0.170-0.264 | 1 |
| Oort Cloud Comets | |||
| 1P/Halley | gas | 0.420.06 | 2 |
| C/1996 B2 Hyatuake | gas | 0.580.20 | 3 |
| C/1995 O1 Hale-Bopp | gas | 0.660.16 | 4 |
| 8P/Tuttle | gas | 0.820.29 | 5 |
| C/2009 P1 Garradd | gas | 0.410.04 | 6 |
| C/2002 T7 LINEAR | gas | 0.500.14 | 7 |
| C/2012 F6 Lemmon | gas | 1.300.32 | 8 |
| C/2014 Q2 Lovejoy | gas | 0.280.08 | 8 |
| C/2007 B3 Lulin | gas | 1.12 | 9 |
| 153P/Ikeya-Zhang | gas | 0.560.06 | 10 |
| Jupiter Family Comets | |||
| 103P/Hartley 2 | gas | 0.320.05 | 11 |
| 67P/Churyumov-Gerasimenko | gas | 1.060.14 | 12 |
| 46P/Wirtanen | gas | 0.320.13 | 13 |
| 45P/Honda-Mrkos-Pajdušáková | gas | 0.40 | 14 |
| Source | Phase | HDO/H2O | Ref. |
| 10-3 | |||
| Clustered Class 0 LYSOs | |||
| NGC 1333 IRAS 4A-NW | gas | 0.540.15 | 1, 2 |
| NGC 1333 IRAS 2A | gas | 0.740.21 | 1 |
| NGC 1333 IRAS 4B | gas | 0.590.26 | 1 |
| IRAS 16293-2422 | gas | 0.920.26 | 1 |
| NGC 1333 SVS 13 | ice | 17 | 3 |
| Clustered Class I LYSOs | |||
| NGC 1333 SVS 12 | ice | 5 | 3 |
| Isolated Class 0 LYSOs | |||
| BHR 71-IRS1 | gas | 1.80.4 | 2 |
| B335 | gas | 6.31.5 | 4 |
| L483 | gas | 4.00.5 | 4 |
| Isolated Class 0/I LYSOs | |||
| L1527 IRS | ice | 4.4 | 5 |
| Isolated Class I LYSOs | |||
| V883 Ori | gas | 2.30.6 | 6 |
| L1551 IRS5 | gas | 2.10.8 | 7 |
| L1489 IRS | ice | 8 | 3 |
| TMR1 | ice | 11 | 3 |
| IMYSOs | |||
| IRAS 05390-0728 | ice | 10 | 8 |
| IRAS 08448-4343 | ice | 10 | 8 |
| HOPS 370 | ice | 4.62.2 | 9 |
| MYSOs | |||
| W3 IRS5 | gas | 1.3 | 10 |
| W33A | gas | 3.0 | 10 |
| AFGL 2591 | gas | 3.3 | 10 |
| NGC 7538 IRS1 | gas | 3.8 | 10 |
| Orion KL Hot Core | gas | 3.0 | 11 |
| NGC 6334 I | gas | 0.210.1 | 12 |
| G34.26+0.15 | gas | 0.35-0.75 | 13 |
| NGC 7538 IRS9 | ice | 8.1-11.4 | 8 |
| GL 2136 | ice | 4 | 8 |
| IRAS 20126 | ice | 2.61.4 | 9 |